You have been appointed financial manager of AB Ltd which is listed on the



Johannesburg Stock Exchange. You need to establish the company’s current cost of



capital and you have been provided with the following information:



Ordinary share capital



There are 240 000 shares in issue at a nominal value of R2.00 each. A dividend of



R2.06 as recently been paid and this dividend amount is expected to continue in the



foreseeable future.



17% preference shares – R120 000



The preference shares dividends for the year end 31 December 2020 have been paid.



Similar preference shares are currently trading at 15% yield to maturity.



10% bonds– R1 280 000



The market interest rate for the same class of bonds as AB Ltd, is 12% per annum.



Additional information



AB Ltd’s beta is 1.57. The return on the market is 20%. The current return on Treasury



bills is 8% per annum. The company tax rate is 28% and the company is in a tax paying



position.



Required:



Calculate the WACC

1
Expert's answer
2021-11-24T14:14:20-0500

WACC calculated by CAPM

"ERi=Rf+b1(Rm - Rf)"

"ERi=8+1.57(20-8)"

"ERi=8+1.57(12)"

"ERi=8+18.84"

"ERi=26.84"%

Therefore, WACC=28.84%